# Symmetry-Resolved Entanglement Entropy from Heat Kernels

### v1

@@ 1 Introduction
Traditional approaches to SREE are primarily based on twist operator correlators in conformal field theory (CFT) Murciano:2020vgh and minimal surface prescriptions in holography Belin:2013uta ; Huang:2025lsy . The heat kernel method, owing to its geometric intuition and adaptability to replica manifolds Solodukhin:2011gn ; Dowker:1977zj ; Fursaev:1994in ; Mann:1997hm ; Frolov:1998ea , offers an alternative approach.
@@ 2.1 Entanglement Entropy and Symmetry Resolution
To compute $Z_{n}(q)$, we introduce a fixed background $U(1)$ gauge field that couples to the conserved current Belin:2013uta . This background field is flat ($dA=0$) everywhere except at the entangling surface $\Sigma$, where it carries a nontrivial Wilson line $\oint_{C}A=-in\mu$.
@@
$Z_{n}(\mu)=\operatorname{Tr}\left[\rho_{A}e^{\mu Q_{A}}\right]^{n},$ (15)
@@ 4.3 SREE for $S^{1}\times H^{d-1}$ Background
$F=-(-1)^{f}\frac{1}{2}\int\frac{ds}{s}\operatorname{Tr}K_{S^{1}\times H^{d-1}}(s)e^{-m_{c}(1-f)s},$ (72)
where $f$ encodes the particle statistics ($f=1$ for Dirac fermions, $f=0$ for scalars) and $m_{c}=-\frac{(d-2)^{2}}{4R^{2}}$ represents the conformal coupling mass with $R=1$.
@@
$K_{S^{1}\times H^{d-1}}(s;x_{1}+2\pi\alpha,\dots,x_{d},y_{1},\dots,y_{d})=(-1)^{f}e^{i\alpha\mu_{E}}K_{S^{1}\times H^{d-1}}(s;x,y),$ (73)
@@
$K_{\mathbb{R}^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}e^{\frac{-(x_{1}-y_{1})^{2}}{4s}},$ (74)
@@
$K_{S^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}\sum_{m\in\mathbb{Z}}e^{-\frac{(y_{1}-x_{1}+2\pi\alpha m)^{2}}{4s}}e^{(-i\alpha\mu_{E}-i\pi f)m}.$ (75)
@@
For the two-dimensional theory at inverse temperature $\beta_{n}=2\pi n$, the finite-$\mu_{E}$ free energy takes the form:
$F_{n}(\mu_{E})=\frac{(-1)^{f}}{2}(2\pi n)V_{H^{1}}\int\frac{ds}{s}\sum_{m}e^{-\frac{\pi^{2}n^{2}m^{2}}{s}}\frac{1}{\sqrt{4\pi s}}e^{(-in\mu_{E}-i\pi f)m}K_{H^{1}}(\rho=0,s),$ (76)
where $K_{H^{1}}(0,t)=\frac{1}{\sqrt{4\pi t}}$, and $y_{1}-x_{1}=0$ was assumed for the trace. The renormalized free energy eliminates the divergent $m=0$ contribution:
$\hat{F}_{n}(\mu_{E})=\frac{(-1)^{f}}{4}\frac{1}{n\pi^{2}}V_{H^{1}}\sum_{m=1}^{\infty}\frac{e^{i(n\mu_{E}+\pi f)m}+e^{-i(n\mu_{E}+\pi f)m}}{m^{2}}.$ (77)
To compute the free energy using the continuous version of the Sommerfeld formula, we employ the auxiliary function:
$G_{m}(d,\mu_{E})\equiv\frac{e^{-im\mu_{E}}}{m^{d}}.$ (78)
For bosons (i.e., $f=0$), the free energy becomes:
$\hat{F}_{n}(\mu_{E})=\frac{1}{4n\pi^{2}}V_{H^{1}}\sum_{m=1}^{\infty}\left[G_{m}(2,n\mu_{E})+G_{m}(2,-n\mu_{E})\right].$ (79)
Consider the function $\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}$, whose poles are located at $\{m\in\mathbb{Z}\}$. By inserting this function to adjust the position of the poles, we can transform the sum into a contour integral:
$\hat{F}_{n}(\mu_{E})$ $=\frac{1}{4n\pi^{2}}V_{H^{1}}\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}).$ (80)
@@
$C_{2}(n,\mu_{E})=\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}).$ (81)
@@
$C_{2}(n,\mu_{E})$ $=-2\pi i\,\mathrm{sgn}(\mu_{E})\,\mathrm{Res}\left(\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(2,n\mu_{E}),m=0\right)$ (82)
$=\frac{1}{2}\mu_{E}^{2}n^{2}-\pi n|\mu_{E}|+\frac{1}{3}\pi^{2}.$ (83)
Consequently, the free energy and Rényi entropy become:
$\hat{F}_{n}(\mu_{E})$ $=\frac{1}{4n\pi^{2}}V_{H^{1}}\left(\frac{1}{2}\mu_{E}^{2}n^{2}-\pi n|\mu_{E}|+\frac{1}{3}\pi^{2}\right),$ (84)
$S_{n}(\mu_{E})$ $=\left[\frac{1}{12}\left(1+\frac{1}{n}\right)-\frac{|\mu_{E}|}{4\pi}\right]V_{H^{1}}.$ (85)
@@
$\hat{F}_{n}(\mu_{E})$ $=\frac{(-1)^{f}}{2}(2\pi n)V_{H^{3}}\int\frac{ds}{s}\sum_{m\in\mathbb{Z}^{+}}e^{-\frac{\pi^{2}n^{2}m^{2}}{s}}\frac{2\cos((n\mu_{E}+\pi f)m)}{\sqrt{4\pi s}}K_{H^{1}}(0,s)$ (86)
$=\frac{(-1)^{f}}{16}\frac{1}{n^{3}\pi^{5}}V_{H^{3}}\sum_{m=1}^{\infty}\frac{e^{i(n\mu_{E}+\pi f)m}+e^{-i(n\mu_{E}+\pi f)m}}{m^{4}}.$ (87)
Applying the Sommerfeld transformation with $G_{m}(4,\mu_{E})\equiv e^{-im\mu_{E}}/m^{4}$, the free energy and coefficient function become:
$\hat{F}_{n}(\mu_{E})=\frac{1}{16n^{3}\pi^{5}}V_{H^{3}}\mathrm{sgn}(\mu_{E})\int_{\Gamma}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(4,n\mu_{E}),$ (88)
$C_{2}(n,\mu_{E})$ $=\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi im\,\mathrm{sgn}(\mu_{E})}}G_{m}(4,n\mu_{E})$ (89)
$=-\frac{1}{24}n^{4}\mu_{E}^{4}+\frac{1}{6}\pi n^{3}|\mu_{E}|^{3}-\frac{1}{6}\pi^{2}n^{2}\mu_{E}^{2}+\frac{1}{45}\pi^{4}.$ (90)
@@
$\hat{F}_{n}(\mu_{E})=\left(\frac{|\mu_{E}|^{3}}{96\pi^{4}}+\frac{1}{720\pi n^{3}}-\frac{\mu_{E}^{4}n}{384\pi^{5}}-\frac{\mu_{E}^{2}}{96\pi^{3}n}\right)V_{H^{3}},$ (91)
$S_{n}(\mu_{E})=\left(\frac{(n+1)\mu_{E}^{2}}{48\pi^{2}n}-\frac{|\mu_{E}|^{3}}{48\pi^{3}}-\frac{1+n+n^{2}+n^{3}}{360n^{3}}\right)\frac{V_{H^{3}}}{2\pi}.$ (92)
These calculations demonstrate the systematic application of our improved Sommerfeld formalism to curved backgrounds, reproducing the expected dimensional scaling while incorporating the effects of finite chemical potential. The results provide explicit validation of our framework in holographically relevant geometries.
@@ Acknowledgements.
We thank Giuseppe Di Giulio and Suting Zhao for helpful discussions. ZYX acknowledges funding from the DFG through the Collaborative Research Center SFB 1170 “ToCoTronics” (Project ID 258499086 – SFB 1170) and support from the Berlin Quantum Initiative.

### v2

@@ 1 Introduction
Traditional approaches to SREE are primarily based on twist operator correlators in conformal field theory (CFT) Murciano:2020vgh and minimal surface prescriptions in holography Belin:2013uta; Huang:2025lsy; zhao2021symmetry; weisenberger2021symmetry; zhao2022charged. Complementary studies have also explored free bosonic systems using correlation matrix techniques Pirmoradian:2023uvt, providing additional examples of how SREE can be computed numerically. The heat kernel method, originating from the seminal works of Fock, Schwinger, and later DeWitt, has become a fundamental tool in quantum field theory and mathematical physics fock1937eigenzeit; schwinger1951gauge; dewitt1965dynamical; vassilevich2003heat; Ivanov:2021fgt. Given its geometric nature and natural compatibility with replica manifold techniques Solodukhin:2011gn; Dowker:1977zj; Fursaev:1994in; Mann:1997hm; Frolov:1998ea, the heat kernel framework provides a promising basis for developing a formulation of SREE.
@@ 2.1 Entanglement Entropy and Symmetry Resolution
To compute $Z_{n}(q)$, we introduce a fixed background $U(1)$ gauge field that couples to the conserved current Belin:2013uta . This background field is flat ($dA=0$) everywhere except at the entangling surface $\Sigma$, where it carries a nontrivial Wilson line.
@@
$Z_{n}(\mu)=\operatorname{Tr}\left[\rho^{n}_{A}e^{i\mu Q_{A}}\right],$ (15)
@@ 4.3 SREE for $S^{1}\times H^{d-1}$ Background
$F=-\frac{1}{2}\int\frac{ds}{s}\operatorname{Tr}2K_{S^{1}\times H^{d-1}}(s),$ (72)
where the factor of 2 in front of the heat kernel appears because a complex scalar field has twice the degrees of freedom of a real scalar field.
@@
$K_{S^{1}\times H^{d-1}}(s;x_{1}+2\pi m\alpha,\dots,x_{d},y_{1},\dots,y_{d})=e^{im\mu_{E}}K_{S^{1}\times H^{d-1}}(s;x,y),$ (73)
@@
$K_{\mathbb{R}^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}e^{\frac{-(y_{1}-x_{1})^{2}}{4s}},$ (74)
@@
$K_{S^{1}}(s;x_{1},y_{1})=\frac{1}{(4\pi s)^{1/2}}\sum_{m\in\mathbb{Z}}e^{-\frac{(y_{1}-x_{1}+2\pi\alpha m)^{2}}{4s}}e^{-im\mu_{E}}.$ (75)
@@
For the two-dimensional theory at inverse temperature $\beta=2\pi\alpha$, the finite-$\mu_{E}$ free energy takes the form:
$F_{\alpha}(\mu_{E})=-\frac{1}{2}(2\pi\alpha)V_{H^{1}{}}\int\frac{ds}{s}\frac{2}{(4\pi s)^{1/2}}\sum_{m\in\mathbb{Z}}e^{-\frac{\pi^{2}\alpha^{2}m^{2}}{s}}e^{-im\mu_{E}}K_{H^{1}}(0,s),$ (76)
where $K_{H^{1}}(0,s)=\frac{1}{\sqrt{4\pi s}}$, and $y_{1}-x_{1}=0$ was assumed for the trace. The renormalized free energy eliminates the divergent $m=0$ contribution:
$\hat{F}_{\alpha}(\mu_{E})=-\frac{V_{H^{1}}}{2\pi^{2}\alpha}\sum_{m\in\mathbb{Z},m\neq 0}\frac{e^{-im\mu_{E}}}{m^{2}}.$ (77)
Consider the function $\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}$, whose poles are located at $\{z=m\in\mathbb{Z}\}$. By inserting this function to adjust the position of the poles, we can transform the sum into a contour integral:
$\hat{F}_{\alpha}(\mu_{E})=-\frac{V_{H^{1}}}{2\pi^{2}\alpha}\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}\frac{e^{-iz\mu_{E}}}{z^{2}}dz.$ (78)
@@
$C_{2}(\mu_{E})=\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}\frac{e^{-iz\mu_{E}}}{z^{2}}dz.$ (79)
@@
$C_{2}(\mu_{E})$ $=-2\pi i\,\mathrm{sgn}(\mu_{E})\,\mathrm{Res}\left(\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}\frac{e^{-iz\mu_{E}}}{z^{2}},z=0\right)$ (80)
$=\frac{1}{2}\mu_{E}^{2}-\pi|\mu_{E}|+\frac{1}{3}\pi^{2}.$ (81)
Consequently, the free energy and entropy become as follows:
$\hat{F}_{\alpha}(\mu_{E})$ $=-\frac{V_{H^{1}}}{12\pi^{2}\alpha}\left(3\mu_{E}^{2}-6\pi|\mu_{E}|+2\pi^{2}\right),$ (82)
$S(\mu_{E})$ $=\left(\frac{1}{3}-\frac{|\mu|}{\pi}+\frac{\mu^{2}}{2\pi^{2}}\right)V_{H^{1}}.$ (83)
@@
$\hat{F}_{\alpha}(\mu_{E})$ $=-\frac{1}{2}(2\pi\alpha)V_{H^{3}}\int\frac{ds}{s}\frac{2}{(4\pi s)^{1/2}}\sum_{m\in\mathbb{Z},m\neq 0}e^{-\frac{\pi^{2}\alpha^{2}m^{2}}{s}}e^{-im\mu_{E}}K_{H^{3}}(0,s)$ (84)
$=-\frac{V_{H^{3}}}{8\alpha^{3}\pi^{5}}\sum_{m\in\mathbb{Z},m\neq 0}\frac{e^{-im\mu_{E}}}{m^{4}}.$ (85)
Applying the Sommerfeld transformation, the free energy and coefficient function become
$\hat{F}_{\alpha}(\mu_{E})=-\frac{V_{H^{3}}}{8\alpha^{3}\pi^{5}}\mathrm{sgn}(\mu_{E})\int_{\Gamma_{2}}\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}\frac{e^{-iz\mu_{E}}}{z^{4}}dz,$ (86)
$C_{2}(\mu_{E})$ $=-2\pi i\,\mathrm{sgn}(\mu_{E})\,\mathrm{Res}\left(\frac{1}{1-e^{-2\pi iz\,\mathrm{sgn}(\mu_{E})}}\frac{e^{-iz\mu_{E}}}{z^{4}},z=0\right)$ (87)
$=\frac{1}{360}\left(8\pi^{4}-60\pi^{2}\mu_{E}^{2}+60\pi|\mu_{E}|^{3}-15\mu_{E}^{4}\right).$ (88)
@@
$\hat{F}_{\alpha}(\mu_{E})=-\frac{V_{H^{3}}}{2880\ \alpha^{3}\pi^{5}}\left(8\pi^{4}-60\pi^{2}\mu_{E}^{2}+60\pi|\mu_{E}|^{3}-15\mu_{E}^{4}\right),$ (89)
$S(\mu_{E})=\left(\frac{1}{45}-\frac{\mu^{2}}{6\pi^{2}}+\frac{|\mu|^{3}}{6\pi^{3}}-\frac{\mu^{4}}{24\pi^{4}}\right)\frac{V_{H^{3}}}{2\pi}.$ (90)
These calculations demonstrate the systematic application of our improved Sommerfeld formalism to curved backgrounds, reproducing the expected dimensional scaling while incorporating the effects of finite chemical potential. Notably, the free energy formulas derived above match exactly the results of Huang:2025lsy, further confirming the correctness of our calculations.
@@ Acknowledgements.
We thank Giuseppe Di Giulio, Aleksandr Ivanov, Mohammad Reza Tanhayi, and Suting Zhao for helpful discussions. ZYX acknowledges funding from the DFG through the Collaborative Research Center SFB 1170 “ToCoTronics” (Project ID 258499086 – SFB 1170) and support from the Berlin Quantum Initiative.